A software firm uses programmers (labor) and computers (capital) to produce apps for mobile devices. The firm estimates that the MP of labor is 5 apps per month and the price of labor is $1000 per month, while the MP of capital is 9 apps per month and its price is $1000. To maximize its profits, what is the rule that the firm uses to determine how much labor and capital to use?
Answer #1

Marginal product of labor is 5 apps

price of labor is $1000

Marginal product of capital is 9 apps.

Price of capital is $1000

At profit maximization point following condition must holds:

(MP of labor / Price of labor) = (MP of capital / Price of capital)

Firm should use that combination of labor and capital which satisfies the above stated condition.

----------

(MP of labor / Price of labor) = (5 /1000) = 0.005

(MP of capital / Price of capital) = (9 / 1000) = 0.009

(MP of capital / Price of capital) is greater than (MP of labor / Price of labor). It implies firm should increase the use of capital and decrease the use of labor till the ratio of marginal product to price of each input equals to each other,